-42=-18-6c
a. c= -24
b. c=144
c. c=4
d. c=10
-42=-18-6c
-42+(18)=-18 +(18)-6c
-24=-6c
-24+6=-6+6
-18=c NOT ONE OF MY ANSWERS

Answer by stanbon(75887)

-42=-18-6c
a. c= -24
b. c=144
c. c=4
d. c=10
-42=-18-6c
-42+(18)=-18 +(18)-6c
-24=-6c
---
At this point you need to divide both sides by -6
to get c = 4
==================
Cheers,
Stan H.
==============